OpenStudy (anonymous):

3a-6 over a+ 2 divided by a-2 over a+2 = ?

zepdrix (zepdrix):

\[\frac{ \frac{ 3a-6 }{ a+2 } }{ \frac{ a-2 }{ a+2 } }\]So is this what the problem looks like?

zepdrix (zepdrix):

Remember how to divide fractions? :) Dividing fractions is the same as Multiplying by the reciprocal of the denominator. (You change it to multiplication and "flip" the bottom fraction). So it would look like this. \[\frac{ 3a-6 }{ a+2 } * \frac{ a+2 }{ a-2 }\] And you can do some nice cancellations from here.. c: post if you're still stuck.